OpenStudy (anonymous):

What is the measure of each interior angle in a regular 30-gon?

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

so the last one is 13?

OpenStudy (wolf1728):

The answer to the PREVIOUS question is 13 sides yes 13 sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my guess is 28 on this one

OpenStudy (wolf1728):

Let's see angle total = sides-2 * 180 in a 30 gon 28*180= 5,040 so EACH angle = 5,040 / 30 = 168

OpenStudy (anonymous):

ok 168 is the answer

OpenStudy (wolf1728):

all right !!

OpenStudy (anonymous):

awesome I got 80% on my test and that's all I need thank you

OpenStudy (wolf1728):

glad to help - did you get the 13 side question right?

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

okay - I'm glad my correction appeared in time okay cya

OpenStudy (anonymous):

bye thanks again

OpenStudy (wolf1728):

u r welcome